OpenStudy (anonymous):

Find a formula for 1/2 + 1/4 + 1/8 + ... + 1/(2^n) by examining the values of this expression for small n. Prove your formula.

OpenStudy (anonymous):

Do by induction

OpenStudy (anonymous):

I can prove it, but I do not know how to find the formula itself.

OpenStudy (anonymous):

The formula is \[\sum_{k=m}^{n} r^k=\frac{r^m-r^{n+1}}{1-r}\] So r=2, m=1 and n remains as n so \[\sum_{k=1}^{n} (\frac{1}{2})^k=\frac{(\frac{1}{2})^1-(\frac{1}{2})^{n+1}}{1-(\frac{1}{2})}\] simplify \[\sum_{k=1}^{n} (\frac{1}{2})^k = 1-(\frac{1}{2})^{n}\] This is a Geometric Series.